OpenStudy (anonymous):

Suppose an isosceles triangle ABC has A = pi/6 and b = c = 5. What is the length of a^2?

OpenStudy (raden):

use the cosine rule : a^2 = b^2 + c^2 - 2bc*cosA

OpenStudy (anonymous):

i never learned that

OpenStudy (anonymous):

these are my options

OpenStudy (raden):

u have learned, now :) just put the value of them ... a^2 = b^2 + c^2 - 2bc*cosA a^2 = 5^2 + 5^2 - 2(5)(5)cos(pi/6) a^2 = 25+25 - 50*1/2*sqrt(3) a^2 = 50 - 25 sqrt(3) a^2 = 25 (2 - sqrt(3)) a^2 = 5^2 (2 - sqrt(3))

OpenStudy (anonymous):

thankyou so much... clear example!!!:)

OpenStudy (raden):

very welcome :)